February 2011

RESOLVED: Network issue

Date: 
Wed, 09/02/2011 - 22:44

We have re-installed the Junos software on one of our firewalls and this is now back online. This took longer than expected due to the fact the partition table on the device was corrupted.

We have applied a backup configuration to the firewall so most customer services are now working but there are a handful of customers whose services will still be unavailable. We are working as quickly as possible to restore service to these customers.

Network issue

Date: 
Wed, 09/02/2011 - 16:42

We are currently experiencing an unexpected issue with one of our firewalls in Docklands, this has caused a loss of connectivity to all customer services. We are currently working on bringing the firewall back online and we will then investigate the root cause.

Please accept our apologies for any inconvenience.

Update 17:10: We have turned off the firewall in the cluster that we believe to be faulty and are currently working on bringing the other firewall back online. This issue is being dealt with at the highest priority, thank you for your patience.

COMPLETE: Virtual Server Migration

Date: 
Sun, 06/02/2011 - 00:00 - 03:00

We will be migrating all our virtual servers at our primary site to new hardware during this window. We don't anticipate more than 60 minutes
of downtime for each virtual server, but we will post updates and any issues on this page.

This will affect VPS, shared and resller hosting services, as well as our own website and billing and support systems. This page will not be affected as it is hosted at our DR site. Most VPS customers will have a change to their package during this window - thse customers have been emailed individually about this.

RESOLVED: Networking issue

Date: 
Sat, 05/02/2011 - 17:16

The issue affecting our switches has now been resolved. Both switches are now back online (and have also been upgraded to the recommended software version).

Customer virtual machines will need to be rebooted before they will be available again (as connectivity to the storage was lost during the switch issue). This is happening now and all servers will be back online shortly.

Please accept our apologies for any inconvenience caused.

Networking issue

Date: 
Sat, 05/02/2011 - 17:04

We are currently experiencing an issue with one switch in our virtual chassis which has caused our storage network to become unavailable, this is having a knock on effect on all services.

We are currently working to get the switch back online and resume normal service. We will post updates here as we know more.

Please accept our apologies for any inconvenience.

XenServer Host Issue

Date: 
Thu, 03/02/2011 - 18:57 - 19:27

One of our XenServer hosts has become unresponsive and we need to restart the host as a result. This will cause a disruption to some customer virtual machines.

We apologise for any inconvenience caused.

Update 19:27